OpenStudy (anonymous):

solve and show steps 4d − 12 = 12 − 4d

OpenStudy (anonymous):

4d+4d=12+12 8d=24 d=24/8 d=3

OpenStudy (anonymous):

no

OpenStudy (anonymous):

4d-12 = 12-4d +12 +12 4d = 24-4d +4d +4d 8d = 24 /8 /8 d=3